# Crosscap states and duality of Ising field theory in two dimensions

### v1

@@ Abstract
where $\mathcal{T}$ stands for time-ordering and $H_{1}(\tau)$ is the perturbation term in the interaction picture, $\langle\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}\rangle_{\mathrm{c}}$ denotes the connected contribution of $\langle\psi_{0}(0)|\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}|\psi_{0}(0)\rangle$, and $\delta c(s)=\lim_{\beta\to\infty}\frac{6L}{\pi\beta}\langle\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}\rangle_{\mathrm{c}}$ is the change of the “running” central charge $c(s)\equiv\delta c(s)+c$. With the knowledge of the CFT correlators (on $\mathbb{RP}^{2}$ and the plane), the expansion of the time-ordered exponential $\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}$ (in powers of the coupling $g$) allows us to calculate the crosscap overlap in a perturbative way. For rational CFTs with relevant perturbations, we have obtained a general result for the first-order correction (if nonvanishing) to the crosscap overlap [43].
The perturbation strategy outlined above is certainly applicable to the Ising field theory. Below we first discuss the crosscap overlaps with $|\mathcal{C}_{+}\rangle$ and then comment on those for $|\mathcal{C}_{-}\rangle$:
(i) For the thermal perturbation [$g_{1}\neq 0$ and $g_{2}=0$ in Eq. (1)], the crosscap overlap $\langle\psi_{0}(s_{1})|\mathcal{C}_{+}\rangle$, with $s_{1}=g_{1}L$, equals the square root of the Klein bottle entropy obtained non-perturbatively in Ref. [40]. Our perturbative expansion agrees with the non-perturbative result order by order [43].
(ii) For the magnetic perturbation [$g_{1}=0$ and $g_{2}\neq 0$ in Eq. (1)], our conformal perturbation calculation indicates that the crosscap overlap should be an even function of the dimensionless coupling: $\langle\psi_{0}(s_{2})|\mathcal{C}_{+}\rangle=\sum_{n=0}^{\infty}\mathcal{C}_{2n}s_{2}^{2n}$, with $s_{2}=g_{2}L^{15/8}$, since crosscap correlators with odd number of $\sigma$-fields vanish. Utilizing the two-point crosscap correlator [Eq. (12)], we obtain the leading-order correction $\mathcal{C}_{2}\approx-1.63528$, which is close to the numerically estimated value $\mathcal{C}_{2}\approx-1.59$ in a lattice model simulation [43].
The dual crosscap overlap $\langle\psi_{0}(s_{\alpha})|\mathcal{C}_{-}\rangle$ is related to $\langle\psi_{0}(s_{\alpha})|\mathcal{C}_{+}\rangle$ ($\alpha=1,2$) through the Kramers-Wannier duality: $\langle\psi_{0}(s_{\alpha})|\mathcal{C}_{-}\rangle=\langle\tilde{\psi}_{0}(s_{\alpha})|\mathcal{C}_{+}\rangle$, where $|\tilde{\psi}_{0}(s_{\alpha})\rangle\equiv U_{\mathrm{KW}}|\psi_{0}(s_{\alpha})\rangle$ denotes the ground state perturbed by the dual operator. For the thermal perturbation, $\varepsilon\leftrightarrow-\varepsilon$ under the Kramer-Wannier duality transformation gives $\langle\psi_{0}(s_{1})|\mathcal{C}_{-}\rangle=\langle\psi_{0}(-s_{1})|\mathcal{C}_{+}\rangle$. For the magnetic perturbation, the dual crosscap overlap $\langle\psi_{0}(s_{2})|\mathcal{C}_{-}\rangle$ differs from $\langle\psi_{0}(s_{2})|\mathcal{C}_{+}\rangle$, as the crosscap correlator of the $\sigma$ field differs from that of the $\mu$ field.
The overlap of the crosscap states $|\mathcal{C}_{\pm}\rangle$ with any perturbed excited states should be a universal scaling function, too. Specifically, for perturbed states that are deformed from primary states, e.g., $|\sigma,\bar{\sigma}\rangle$ and $|\varepsilon,\bar{\varepsilon}\rangle$ in the Ising CFT, a similar conformal perturbation approach can be developed. This involves a slight modification to Eq. (15) by inserting the corresponding primary field at infinity ($\tau=\infty$), in accordance with the state-operator correspondence in CFT [55]. A concrete example confirming that excited-state crosscap overlaps are also universal scaling functions is the 2D Ising CFT with thermal perturbation. As the thermal perturbation takes a quadratic form in the fermionic representation, we find the following universal overlaps between $|\mathcal{C}_{\pm}\rangle$ [Eq. (10)] and perturbed eigenstates [43]:
$|\langle\psi_{n_{1}\cdots n_{M}}(s_{1})|\mathcal{C}_{\pm}\rangle|^{2}=1+\frac{(-1)^{M}}{\sqrt{1+e^{\mp 2\pi s_{1}}}}\,,$ (16)
where $|\psi_{n_{1}\cdots n_{M}}(s_{1})\rangle$ is the eigenstate deformed from the unperturbed state $|\psi_{n_{1}\cdots n_{M}}(0)\rangle\equiv\prod_{\alpha=1}^{M}b^{\dagger}_{n_{\alpha}-1/2}\bar{b}^{\dagger}_{n_{\alpha}-1/2}|0\rangle_{\mathrm{NS}}$ in the Ising CFT basis.
@@
Summary and outlook — In conclusion, two distinct crosscap states in the 2D Ising field theory that are connected to each other via the Kramers-Wannier duality have been thoroughly investigated. The Majorana free field representation of these crosscap states are established and the bosonization approach is extended to calculate the crosscap correlators, i.e., correlation functions on the real projective plane ($\mathbb{RP}^{2}$). To understand the crosscap overlap in the vicinity of a critical point, we have developed a conformal perturbation theory to compute the low order coefficients in the power series expansion of the overlap. These results provide vital insights about field theories on non-orientable manifolds and also enrich the numerical toolbox for identifying critical theories in lattice models (e.g., Ref. [56]).
Apart from the 2D Ising field theory, duality has also been uncovered in many other field theories (e.g., $\mathbb{Z}_{N}$ parafermion CFTs [57]). It would be interesting to explore new crosscap states enriched by duality. This may shed light on a conjectured link between the Klein bottle entropy and the renormalization group flow [40], in analogous to the $c$-theorem [58] for central charge and the $g$-theorem [59, 60] for Affleck-Ludwig boundary entropy. Moreover, crosscap overlaps on lattices provide a promising route for extracting crosscap coefficients of 3D CFTs, which would complement the bootstrap approach where only ratios of crosscap coefficients were obtained [32].
@@
This Supplemental Material provides derivation details of some results in the main text. In Sec. I, we briefly review the exact solution of the transverse field Ising chain (TFIC) and its Kramers-Wannier duality. In Sec. II, we introduce the lattice crosscap states and show how they are related by the Kramers-Wannier duality transformation. In Sec. III, we derive the fermionic representation of the lattice crosscap states and calculate their overlaps with the eigenstates of the TFIC. In Sec. IV, we identify the fermionic representation of the conformal crosscap states in the Ising conformal field theory (CFT). The exact crosscap overlaps are derived for the Ising CFT in the presence of thermal perturbation. In Sec. V, we discuss the bosonization of the crosscap states and use it to calculate the crosscap correlators in the Ising CFT. In Sec. VI, we develop the conformal perturbation theory to calculate the crosscap overlaps as universal scaling functions.
@@
where we consider even $N$ and $h>0$. The Hamiltonian has $\mathbb{Z}_{2}$ symmetry, $[H_{\mathrm{latt}},Q]=0$ with $Q=\prod_{j=1}^{N}\sigma_{j}^{z}$. In the following, we focus on the $\mathbb{Z}_{2}$ even ($Q=1$) subspace, which is called the Neveu-Schwarz (NS) sector following the CFT convention.
@@
[…]
@@
$H_{\mathrm{latt}}=-\frac{i}{2}\sum_{j=1}^{N}(\chi_{j}-\bar{\chi}_{j})(\chi_{j+1}+\bar{\chi}_{j+1})-ih\sum_{j=1}^{N}\chi_{j}\bar{\chi}_{j}\,,$ (S8)
@@
$H_{\mathrm{latt}}=-i\sum_{j=1}^{N}\bar{\chi}^{\prime}_{j}\chi^{\prime}_{j+1}-ih\sum_{j=1}^{N}\chi^{\prime}_{j}\bar{\chi}^{\prime}_{j}\,.$ (S11)
@@
$U_{\mathrm{KW}}\bar{\chi}^{\prime}_{j}U_{\mathrm{KW}}^{\dagger}$ =$=\begin{cases}\chi^{\prime}_{j+1}&\quad j=1,2,\ldots,N-1\\ -\chi^{\prime}_{1}&\quad j=N\\ \end{cases}\,.~{}$ (S13)
@@
[…]
@@
$\sigma_{j}^{z}$ $=-i\chi^{\prime}_{j}\bar{\chi}^{\prime}_{j}\,,\quad\sigma^{x}_{j}=(-)^{j}\prod_{l=1}^{j-1}(i\chi^{\prime}_{l}\bar{\chi}^{\prime}_{l})\chi^{\prime}_{j}\,,$
@@
$U_{\mathrm{KW}}|\psi_{0}(h)\rangle=|\psi_{0}(1/h)\rangle\,,$ (S18)
@@
In this section, we consider the lattice crosscap states and discuss how they are related via the Kramers-Wannier duality.
The lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$ is the product of Bell pair states identifying Ising spins at two antipodal sites:
@@
$U_{\mathrm{KW}}\sigma_{j}^{x}U_{\mathrm{KW}}^{\dagger}$ $=-\prod_{l=1}^{j-1}\left[U_{\mathrm{KW}}\left(-i\chi^{\prime}_{l}\bar{\chi}^{\prime}_{l}\right)U_{\mathrm{KW}}^{\dagger}\right](U_{\mathrm{KW}}\chi^{\prime}_{j}U_{\mathrm{KW}}^{\dagger})$
$=-\prod_{l=1}^{j-1}\left(-i\bar{\chi}^{\prime}_{l}\chi^{\prime}_{l+1}\right)\bar{\chi}^{\prime}_{j}$
$=-\bar{\chi}^{\prime}_{1}\prod_{l=2}^{j}(-i\chi^{\prime}_{l}\bar{\chi}^{\prime}_{l})$
$=\sigma_{1}^{y}\prod_{l=2}^{j}\sigma_{l}^{z}\,.$ (S31)
Since the lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$ lives in the NS sector, $Q|\mathcal{C}_{\mathrm{latt}}^{+}\rangle=|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$, it is natural to consider its Kramers-Wannier dual state, denotes as $|\mathcal{C}_{\mathrm{latt}}^{-}\rangle\equiv U_{\mathrm{KW}}|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$, which also lives in the NS sector. The explicit form of the (dual) lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}^{-}\rangle$ reads
$|\mathcal{C}_{\mathrm{latt}}^{-}\rangle\equiv U_{\mathrm{KW}}|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$ $=\prod_{j=1}^{N/2}\left(1+U_{\mathrm{KW}}\sigma_{j}^{x}U_{\mathrm{KW}}^{\dagger}U_{\mathrm{KW}}\sigma_{j+1}^{x}U_{\mathrm{KW}}^{\dagger}\right)\left(U_{\mathrm{KW}}|\Uparrow\rangle\right)$
$=\prod_{j=1}^{N/2}\left(1+\mu_{j}\mu_{j+N/2}\right)\frac{1}{\sqrt{2}}\left(|\Rightarrow\rangle+|\Leftarrow\rangle\right)\,,~{}$ (S32)
@@
In this section, we calculate the overlaps of the lattice crosscap states $|\mathcal{C}_{\mathrm{latt}}^{\pm}\rangle$ with eigenstates of the TFIC.
We first derive the fermionic representation of the lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$, which is a superposition of two fermionic Gaussian states in the NS sector. This crucial observation allows us to calculate its overlap with the ground state $|\psi_{0}(h)\rangle$ of the TFIC as well as all excited states. The crosscap overlap results for the (dual) lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}^{-}\rangle$ are obtained via the Kramers-Wannier duality.
C.1 A. Fermionic representation of $|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$

### v2

@@ Abstract
where $\mathcal{T}$ stands for time-ordering and $H_{1}(\tau)$ is the perturbation term in the interaction picture, $\langle\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}\rangle_{\mathrm{c}}$ denotes the connected contribution of $\langle\psi_{0}(0)|\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}|\psi_{0}(0)\rangle$, and $\delta c(s)=\lim_{\beta\to\infty}\frac{6L}{\pi\beta}\langle\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}\rangle_{\mathrm{c}}$ is the change of the “running” central charge $c(s)\equiv\delta c(s)+c$. With the knowledge of the CFT correlators (on $\mathbb{RP}^{2}$ and the plane), the expansion of the time-ordered exponential $\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}$ (in powers of the coupling $g$) allows us to calculate the crosscap overlap in a perturbative way.
We obtain a general expression for the first-order correction to the crosscap overlap [46]:
$\delta\langle\psi_{0}(s)|\mathcal{C}\rangle$ $=\mathcal{A}\cdot\frac{{}_{2}F_{1}(h,2h,h+1;-1)}{2(2\pi)^{1-2h}\,h}\cdot s+\cdots\,$, (18)
where $\mathcal{A}=\lim_{w\to 0}\langle\psi_{0}(0)|\varphi(w)|\mathcal{C}\rangle$ is the crosscap coefficient in the one-point conformal crosscap correlator, and $_{2}F_{1}(a,b,c;x)$ is the hypergeometric function. Since $_{2}F_{1}(h,2h,h+1;-1)>0$ for $h<1$, the sign of the first-order correction is solely determined by $\mathcal{A}$. As a concrete example, consider the $A$-series unitary minimal models [12], with central charge $c=1-\frac{6}{m(m+1)}$, $m\geq 3$. The primary fields are labeled by integer pairs $(r,s)$ with $1\leq r\leq m-1$ and $1\leq s\leq m$, subject to $r+s=\mathrm{even}$. The crosscap state associated with the identity field $(1,1)$ can be expressed in terms of the Ishibashi states as
$|\mathcal{C}_{(1,1)}\rangle=\sum_{1\leq r\leq m-1\,,1\leq s\leq m}^{r+s=\mathrm{even}}\mathcal{A}_{(r,s)}|(r,s)\rangle\rangle_{\mathcal{C}}\,$, (19)
where $\mathcal{A}_{(r,s)}$ are the crosscap coefficients of the primary fields $\varphi_{(r,s)}$ [57, 23]:
$\mathcal{A}_{(r,s)}=O_{r}O_{s}\sqrt{\frac{\sqrt{2}\cot\left(\frac{\pi r}{2m}\right)\cot\left(\frac{\pi s}{2(m+1)}\right)}{\sqrt{m(m+1)}}}$ (20)
with $O_{n}=\frac{1-(-1)^{n}}{2}$. Since $\mathcal{A}_{(r,s)}\geq 0$ for all $(r,s)$, Eq. (18) provides a perturbative proof of the monotonicity conjecture [43] for the crosscap overlap (and hence Klein bottle entropy) under the renormalization group flow in $A$-series unitary minimal models whenever the crosscap coefficient of the perturbing field is non-vanishing.
The conformal perturbation theory outlined above is directly applicable to the Ising field theory in Eq. (1), where $|\mathcal{C}_{(1,1)}\rangle$ in Eq. (19) reduces to $|\mathcal{C}_{+}\rangle$. The results are summarized as follows: (i) For the thermal perturbation, the crosscap overlap $\langle\psi_{0}(s_{1})|\mathcal{C}_{+}\rangle$, with $s_{1}=g_{1}L$, has been computed non-perturbatively in Ref. [43]. Our conformal perturbation theory calculations agree with the non-perturbative result up to second order [46]. (ii) For the magnetic perturbation, conformal perturbation theory shows that the crosscap overlap is an even function of the dimensionless coupling: $\langle\psi_{0}(s_{2})|\mathcal{C}_{+}\rangle=\sum_{n=0}^{\infty}\mathcal{C}_{2n}s_{2}^{2n}$, with $s_{2}=g_{2}L^{15/8}$, since crosscap correlators with odd number of $\sigma$-fields vanish. Utilizing the two-point crosscap correlator [Eq. (14)], we obtain the leading-order correction $\mathcal{C}_{2}\approx-1.63528$ [46]. This shows that the crosscap overlap reaches its maximum at $s_{2}=0$, thereby supporting the monotonicity conjecture in Ref. [43]. An important future direction is to extend this perturbative proof to other perturbed CFTs in which the first-order correction of the crosscap overlap vanishes.
@@
Summary and outlook — In conclusion, we have thoroughly investigated two distinct crosscap states in the 2D Ising field theory, connected by Kramers-Wannier duality. By developing a conformal perturbation theory for the crosscap overlap, we extended its definition beyond CFT and established a perturbative expansion in the coupling constant. This framework not only deepens our understanding of field theories on non-orientable manifolds, but also enriches the numerical toolbox for identifying critical theories in lattice models [58, 59, 51, 60].
Beyond the Ising model, many field theories exhibit dualities (e.g., $\mathbb{Z}_{N}$ parafermion CFTs [61]), and it is natural to ask how crosscap states transform under such dualities. Exploring this interplay may reveal new structures of non-orientable CFTs. A second intriguing direction is to clarify the link between Klein bottle entropy and renormalization-group flow [43], in analogy with the $c$-theorem [62] and $g$-theorem [63, 64]. Finally, crosscap overlaps may provide a direct route to extract crosscap coefficients in 3D CFTs [60], thereby complementing bootstrap results where only ratios of such coefficients are accessible [33].
@@
This Supplemental Material provides derivation details of some results in the main text. In Sec. I, we briefly review the exact solution of the transverse field Ising chain (TFIC) and its Kramers-Wannier duality. In Sec. II, we introduce the lattice crosscap states and show how they are related by the Kramers-Wannier duality transformation. In Sec. III, we derive the fermionic representation of the lattice crosscap states and calculate their overlaps with the eigenstates of the TFIC. In Sec. IV, we identify the continuum counterparts of the lattices crosscap states and relate them to the Pradisi-Sagnotti-Stanev crosscap states in the Ising conformal field theory (CFT). The exact crosscap overlaps are derived for the Ising CFT in the presence of thermal perturbation. In Sec. V, we discuss the bosonization of the conformal crosscap states and use it to calculate the crosscap correlators in the Ising CFT. In Sec. VI, we develop the conformal perturbation theory to calculate the crosscap overlaps as universal scaling functions. We use it to provide a perturbative proof for the monotonicity of the Klein bottle entropy for the perturbed $A$-series unitary minimal models. We also apply the conformal perturbation theory to calculate the leading-order corrections to the crosscap overlaps in the Ising field theory and the $\mathbb{Z}_{3}$ parafermion CFT with thermal perturbation.
@@
where we consider even $N$ and $h>0$. The Hamiltonian has $\mathbb{Z}_{2}$ symmetry, $H_{\mathrm{latt}},Q]=0$ with $Q=\prod_{j=1}^{N}\sigma_{j}^{z}$. In the following, we focus on the $\mathbb{Z}_{2}$ even ($Q=1$) subspace, which is called the Neveu-Schwarz (NS) sector following the CFT convention.
@@
where $|0\rangle_{c}$, the vacuum of the Jordan-Wigner fermions, satisfies $c_{j}|0\rangle_{c}=0\;\forall j$ and is the fully polarized state in the original $\sigma^{z}$-basis: $|0\rangle_{c}=|\Downarrow\rangle\equiv|\downarrow_{1}\downarrow_{2}\cdots\downarrow_{N}\rangle$. To fix the phase of the ground state, we require that its overlap with $|\Downarrow\rangle$ is positive: $\langle\Downarrow|\psi_{0}(h)\rangle=\prod_{k>0}\sin\frac{\theta_{k}}{2}>0$, and $|\psi_{0}(h)\rangle$ is then a real positive wave function in the $\sigma^{z}$-basis – this agrees with the Perron-Frobenius theorem, as the off-diagonal matrix elements of the Hamiltonian (S1) are non-positive in the $\sigma^{z}$-basis. All excited states in the NS sector are obtained by acting an even number of $d^{{\dagger}}_{k}$’s (with distinct momenta) on top of the ground state $|\psi_{0}(h)\rangle$.
@@
$H_{\mathrm{latt}}=\frac{i}{2}\sum_{j=1}^{N}(\chi_{j}-\bar{\chi}_{j})(\chi_{j+1}+\bar{\chi}_{j+1})-ih\sum_{j=1}^{N}\chi_{j}\bar{\chi}_{j}\,$, (S8)
@@
$H_{\mathrm{latt}}=i\sum_{j=1}^{N}\bar{\chi}^{\prime}_{j}\chi^{\prime}_{j+1}+ih\sum_{j=1}^{N}\chi^{\prime}_{j}\bar{\chi}^{\prime}_{j}\,.$ (S11)
@@
$U_{\mathrm{KW}}\bar{\chi}^{\prime}_{j}U_{\mathrm{KW}}^{\dagger}$ $=\begin{cases}\chi^{\prime}_{j+1}&\quad j=1,2,\ldots,N-1\,,\\ -\chi^{\prime}_{1}&\quad j=N\,.\end{cases}$ (S13)
@@
which transforms the TFIC with transverse field $h$ to the same Hamiltonian but with transverse field $1/h$.
@@
$\sigma_{j}^{z}$ $=-i\chi^{\prime}_{j}\bar{\chi}^{\prime}_{j}\,,\quad\sigma^{x}_{j}=(-1)^{j}\prod_{l=1}^{j-1}(i\chi^{\prime}_{l}\bar{\chi}^{\prime}_{l})\chi^{\prime}_{j}\,$,
@@
$U_{\mathrm{KW}}|\psi_{0}(h)\rangle=|\psi_{0}(1/h)\rangle\,.$ (S18)
@@
In this section, we consider the lattice crosscap states and discuss how they are related via the Kramers-Wannier duality. We also derive an alternative representation of the Kramers-Wannier dual lattice crosscap state, which is expressed as an equal weight superposition of the original lattice crosscap state and another modified lattice crosscap state.
B.1 A. Lattice crosscap state and its dual
The lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$ is the product of Bell pairs identifying Ising spins at two antipodal sites:
@@
$U_{\mathrm{KW}}\sigma_{j}^{x}U_{\mathrm{KW}}^{\dagger}=-\prod_{l=1}^{j-1}\left[U_{\mathrm{KW}}\left(-i\chi^{\prime}_{l}\bar{\chi}^{\prime}_{l}\right)U_{\mathrm{KW}}^{\dagger}\right](U_{\mathrm{KW}}\chi^{\prime}_{j}U_{\mathrm{KW}}^{\dagger})=-\bar{\chi}^{\prime}_{1}\prod_{l=2}^{j}(-i\chi^{\prime}_{l}\bar{\chi}^{\prime}_{l})=\sigma_{1}^{y}\prod_{l=2}^{j}\sigma_{l}^{z}\,.$ (S31)
Since the lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$ lives in the NS sector, $Q|\mathcal{C}_{\mathrm{latt}}^{+}\rangle=|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$, it is natural to consider its Kramers-Wannier dual state, denoted as $|\mathcal{C}_{\mathrm{latt}}^{-}\rangle\equiv U_{\mathrm{KW}}|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$, which also lives in the NS sector. The explicit form of the dual lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}^{-}\rangle$ reads
$|\mathcal{C}_{\mathrm{latt}}^{-}\rangle\equiv U_{\mathrm{KW}}|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$ $=\prod_{j=1}^{N/2}\left(1+U_{\mathrm{KW}}\sigma_{j}^{x}U_{\mathrm{KW}}^{\dagger}\cdot U_{\mathrm{KW}}\sigma_{j+1}^{x}U_{\mathrm{KW}}^{\dagger}\right)\left(U_{\mathrm{KW}}|\Uparrow\rangle\right)$
$=\prod_{j=1}^{N/2}\left(1+\mu_{j}\mu_{j+N/2}\right)\frac{1}{\sqrt{2}}\left(|\Rightarrow\rangle+|\Leftarrow\rangle\right)\,$, (S32)
@@
Moreover, since $U_{\mathrm{KW}}\mu_{j}\mu_{j+N/2}U^{\dagger}_{\mathrm{KW}}=\sigma^{x}_{j+1}\sigma^{x}_{j+1+N/2}$ and $U_{\mathrm{KW}}\,\frac{1}{\sqrt{2}}\left(|\Rightarrow\rangle+|\Leftarrow\rangle\right)=|\Uparrow\rangle$, imposing periodic boundary conditions one finds
$U_{\mathrm{KW}}|\mathcal{C}_{\mathrm{latt}}^{-}\rangle=|\mathcal{C}_{\mathrm{latt}}^{+}\rangle\,.$ (S33)
Thus, the two lattice crosscap states, $|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$ and $|\mathcal{C}_{\mathrm{latt}}^{-}\rangle$, are dual to each other under the Kramers-Wannier transformation, as shown in Eqs. (S32) and  (S33).
B.2 B. Alternative representation of the dual lattice crosscap state
In the main text [Eq. (5)], we mentioned that the dual lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}^{-}\rangle$ can be expressed as
$|\mathcal{C}^{-}_{\mathrm{latt}}\rangle=\frac{1}{\sqrt{2}}\left(|\mathcal{C}^{+}_{\mathrm{latt}}\rangle+|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle\right)\,.$ (S34)
Here, $|\mathcal{C}^{+}_{\mathrm{latt}}\rangle$ is the original lattice corsscap state [Eq. (S28)], and $|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle$ is a modified lattice crosscap state (with a different type of Bell pairs between antipodal sites):
$|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle=\prod_{j=1}^{N/2}\left(|\uparrow_{j}\uparrow_{j+N/2}\rangle-|\downarrow_{j}\downarrow_{j+N/2}\rangle\right)\,.$ (S35)
To prove Eq. (S34), we compare both sides explicitly in the $\sigma^{z}$-basis. We start with the definition of $|\mathcal{C}^{-}_{\mathrm{latt}}\rangle$ in Eq. (S32) and decompose it as
$|\mathcal{C}^{-}_{\mathrm{latt}}\rangle=\frac{1}{\sqrt{2}}\left(|D_{+}\rangle+|D_{-}\rangle\right)$ (S36)
with
$|D_{\pm}\rangle=\frac{1}{2^{N/2}}\prod_{j=1}^{N/2}\left(1+\prod_{l=j+1}^{j+N/2}\sigma_{l}^{z}\right)\prod_{j=1}^{N}(1\pm\sigma_{j}^{x})|\Uparrow\rangle\,.$ (S37)
Since $\sigma^{z}|s\rangle=(-1)^{\frac{s-1}{2}}|s\rangle$ for $s=\pm 1$, one finds
$|D_{\pm}\rangle$ $=\frac{1}{2^{N/2}}\sum_{s_{1},\cdots,s_{N}=\pm 1}\prod_{j=1}^{N/2}\left[1+\prod_{l=j+1}^{j+N/2}(-1)^{\frac{s_{l}-1}{2}}\right](\pm 1)^{\sum_{j=1}^{N}\frac{s_{j}-1}{2}}|s_{1}\cdots s_{N}\rangle$
$=\sum_{s_{1},\cdots,s_{N}=\pm 1}\prod_{j=1}^{N/2}\delta^{(2)}_{\sum_{l=j+1}^{j+N/2}\frac{s_{l}-1}{2},0}\,(\pm 1)^{\sum_{j=1}^{N}\frac{s_{j}-1}{2}}|s_{1}\cdots s_{N}\rangle\,$, (S38)
where $\delta^{(2)}_{n,0}\equiv\frac{1+(-1)^{n}}{2}$ ($n\in\mathbb{Z}$) is the Kronecker-$\delta$ function with period $2$. The relation
$\delta^{(2)}_{\sum_{l=j}^{j-1+N/2}\frac{s_{l}-1}{2},0}\,\delta^{(2)}_{\sum_{l=j+1}^{j+N/2}\frac{s_{l}-1}{2},0}=\delta^{(2)}_{\sum_{l=j}^{j-1+N/2}\frac{s_{l}-1}{2},0}\delta_{s_{j},s_{j+N/2}}\,,\quad j=2,\cdots,N/2\,$ (S39)
allows us to simplify the expression of $|D_{\pm}\rangle$,
$|D_{\pm}\rangle$ $=\sum_{s_{1},\cdots,s_{N}=\pm 1}\delta^{(2)}_{\sum_{j=2}^{N/2+1}\frac{s_{j}-1}{2},0}\,\prod_{j=2}^{N/2}\delta_{s_{j},s_{j+N/2}}(\pm 1)^{\sum_{j=1}^{N}\frac{s_{j}-1}{2}}|s_{1}\cdots s_{N}\rangle$
$=\sum_{s_{1},\cdots,s_{N/2+1}=\pm 1}\delta^{(2)}_{\sum_{j=2}^{N/2+1}\frac{s_{j}-1}{2},0}\,(\pm 1)^{\frac{s_{1}-1}{2}-\frac{s_{N/2+1}-1}{2}+2\sum_{j=2}^{N/2+1}\frac{s_{j}-1}{2}}|s_{1}s_{2}\cdots s_{N/2}s_{N/2+1}s_{2}\cdots s_{N/2}\rangle$
$=\sum_{s_{1},\cdots,s_{N/2+1}=\pm 1}\delta^{(2)}_{\sum_{j=2}^{N/2+1}\frac{s_{j}-1}{2},0}\,(\pm 1)^{\frac{1}{2}(s_{1}-s_{N/2+1})}|s_{1}s_{2}\cdots s_{N/2}s_{N/2+1}s_{2}\cdots s_{N/2}\rangle\,.$ (S40)
Summing $|D_{+}\rangle$ and $|D_{-}\rangle$ gives
$|D_{+}\rangle+|D_{-}\rangle$ $=\sum_{s_{1},\cdots,s_{N/2+1}=\pm 1}\delta^{(2)}_{\sum_{j=2}^{N/2+1}\frac{s_{j}-1}{2},0}\,\left[1+(-1)^{\frac{1}{2}(s_{1}-s_{N/2+1})}\right]|s_{1}s_{2}\cdots s_{N/2}s_{N/2+1}s_{2}\cdots s_{N/2}\rangle$
$=2\sum_{s_{1},\cdots,s_{N/2+1}=\pm 1}\delta^{(2)}_{\sum_{j=2}^{N/2+1}\frac{s_{j}-1}{2},0}\delta_{s_{1},s_{N/2+1}}|s_{1}s_{2}\cdots s_{N/2}s_{N/2+1}s_{2}\cdots s_{N/2}\rangle$
$=2\sum_{s_{1},\cdots,s_{N/2}=\pm 1}\delta^{(2)}_{\sum_{j=1}^{N/2}\frac{s_{j}-1}{2},0}|s_{1}\cdots s_{N/2}s_{1}\cdots s_{N/2}\rangle\,.$ (S41)
On the other hand, the right-hand side of Eq. (S34) expands as
$|\mathcal{C}^{+}_{\mathrm{latt}}\rangle+|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle$ $=\sum_{s_{1},\cdots,s_{N/2}=\pm 1}\left[1+(-1)^{\sum_{j=1}^{N/2}(s_{j}-1)/2}\right]|s_{1}\cdots s_{N/2}s_{1}\cdots s_{N/2}\rangle$
$=2\sum_{s_{1},\cdots,s_{N/2}=\pm 1}\delta^{(2)}_{\sum_{j=1}^{N/2}\frac{s_{j}-1}{2},0}\,|s_{1}\cdots s_{N/2}s_{1}\cdots s_{N/2}\rangle\,$, (S42)
Since the two expansions coincide, Eq. (S34) is proved.
@@
In this section, we calculate the overlaps of the lattice crosscap states $|\mathcal{C}_{\mathrm{latt}}^{\pm}\rangle$, and the modified crosscap state $|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle$, with eigenstates of the TFIC.
We first derive the fermionic representation of the lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}^{+}\rangle$, which is a superposition of two fermionic Gaussian states in the NS sector. This crucial observation allows us to calculate its overlap with the ground state $|\psi_{0}(h)\rangle$ of the TFIC as well as all excited states. The overlaps for the dual state $|\mathcal{C}_{\mathrm{latt}}^{-}\rangle$ then follow directly from Kramers-Wannier duality. Moreover, the modified lattice crosscap state $|\mathcal{C}^{\prime}_{\mathrm{latt}}\rangle$ admits an analogous fermionic representation, allowing its overlaps to be computed as well. The expansions of all three lattice crosscap states in the TFIC eigenbasis provide an alternative verification of Eq. (S34).
C.1 A. Fermionic representation of the lattice crosscap state